# Does VLM Classification Benefit from LLM Description Semantics?

### v1

@@ Abstract
Accurately describing images via text is a foundation of explainable AI. Vision-Language Models (VLMs) like CLIP have recently addressed this by aligning images and texts in a shared embedding space, expressing semantic similarities between vision and language embeddings. VLM classification can be improved with descriptions generated by Large Language Models (LLMs). However, it is difficult to determine the contribution of actual description semantics, as the performance gain may also stem from a semantic-agnostic ensembling effect. Considering this, we ask how to distinguish the actual discriminative power of descriptions from performance boosts that potentially rely on an ensembling effect. To study this, we propose an alternative evaluation scenario that shows a characteristic behavior if the used descriptions have discriminative power. Furthermore, we propose a training-free method to select discriminative descriptions that work independently of classname ensembling effects. The training-free method works in the following way: A test image has a local CLIP label neighborhood, i.e., its top-$k$ label predictions. Then, w.r.t. to a small selection set, we extract descriptions that distinguish each class well in the local neighborhood. Using the selected descriptions, we demonstrate improved classification accuracy across seven datasets and provide in-depth analysis and insights into the explainability of description-based image classification by VLMs.
@@
Figure 1: Are the extra semantics provided by LLM truly useful? Our method first identifies candidate labels using only the class name. We then filter out descriptions that may seem logical but do not differentiate the group, such as those ambiguous, overly generic, or act as noise. This refinement ensures that the remaining descriptions provide distinctive vision-language cues within the local candidate neighborhood, offering more specificity than the class name alone can capture.
@@ 1 Introduction
Human visual recognition is closely tied to verbal reasoning, as it often relies on the ability to express visual information in words (Zhao et al. 2022). However, a neural network usually does not exhibit this property, making its explainability a significant concern for the machine learning community. Some studies (Zhang et al. 2024; Hakimov and Schlangen 2023) aim to connect visual cues and textual descriptions, but these usually require extensive human subject analysis and highly specific datasets with annotations  (Young et al. 2014), which are costly to obtain (Lin et al. 2014).
@@
Vision-Language Models (VLM) (Radford et al. 2021; Li et al. 2021, 2019) tackle this issue by training neural networks to link images and their textual descriptions within a shared embedding space. This enhances the correlation between visual and textual details. VLMs can be applied to zero-shot image classification by passing an image through the VLM’s image encoder and prompting the text encoder with hand-crafted inputs like “a photo of a [classname]” (Radford et al. 2021). Recent work(Menon and Vondrick 2022) extends this approach by incorporating additional descriptions generated by Large Language Models (LLMs) for each class name. LLMs like GPT-3 (Brown et al. 2020; Ouyang et al. 2022) or Llama (Touvron et al. 2023), trained on extensive text corpora, are intended to provide richer semantics, enhancing VLM classification.
@@
However, Does VLM classification truly benefit from LLM-generated description semantics? This work explores this core question as LLM-generated descriptions present several challenges. For example, descriptions can overlap for similar classes — such as parrots and sparrows — both described as having feathers, which is not a distinguishing feature. Moreover, while supplying the model with as many LLM-generated descriptions as possible may seem advantageous, it results in excessively lengthy collections. This complicates understanding the contribution of each description to the final decision.
@@
Another problem is the structured noise ensembling phenomenon (Roth et al. 2023): LLM-generated descriptions can be replaced with random characters and high-level concepts (such as “Baklava”, “a food that is 34mfqr5”) while still improving the model performance. This raises the question of whether the improvement is due to additional semantics of the LLM-generated descriptions or to the ensemble effect of the noise augmentation.
@@
Given these challenges, a proposed model should meet three criteria: 1) As humans who describe with a limited set of descriptions, the model should also operate with a manageable number of text descriptions. 2) These descriptions should be semantically meaningful. 3) The model should be resilient to noise ensembling. To address the issue of structured noise ensembling, we constrain the model to use only textual descriptions that do not contain the classname, i.e. classname-free descriptions.
@@ 3 Method
First, we introduce the conventional task formulation for image classification using Vision-Language Models. We then present our unique approach to this task to enable interpretability. Finally, we propose a specific solution to enhance the results further.
@@ 3.1 Background
where $e(\cdot)$ is the image or language embedding. Another way to obtain vision-language scores is via ensembling. The motivation for that ensembling is from how a human describes an object. For example, when describing an apple, we can describe it as a “green stuff”, “a round object”, or “fruit of the same size as an orange”. In this case, there is no single text representation $t_{c}$ for a class $c$ but a set of language representations $\mathcal{D}(c)$ where the ensembling happens over the elements of $\mathcal{D}(c)$:
@@
Figure 2: Our Method. In the conventional setup (left), using CLIP with LLM-generated class descriptions or random strings can sometimes result in performance gains due to the added semantics or the ensemble effect. However, when the classname is removed, i.e. under the proposed classname-free setup (right), these descriptions will fail to perform well, as only meaningful descriptions w.r.t. the class are useful where random strings or not informative descriptions bring no gain.
@@ 3.2 Our Approach
In the conventional setup (Menon and Vondrick 2022; Pratt et al. 2023; Esfandiarpoor and Bach 2024), performance gains may result from the noise augmentation of the class name text [cls] embedding through its various combinations with [LLM-generated category], [LLM-generated descriptions], and even random strings. While random strings should not contribute extra semantics and be embedded far away from [cls], this can sometimes apply to LLM-assigned text as well due to the vocabulary discrepancy between VLMs and LLMs. Another cause may also be the fact that the images do not exhibit the described property. Despite this, such combinations can still achieve good performance even though the assigned descriptions are not semantically correct.
To better investigate whether the improved performance stems from semantic enrichment or merely from the ensemble effect, we propose an approach where, out of all elements in $\mathcal{D}(c)$, exactly one element should contain the class name $c$. The remaining elements must contain textual descriptions without the class name. This set of descriptions then becomes:
@@
As discussed previously, under the conventional scenario, the evaluation of the language-ensembling VLM method is a simple averaging of aggregated class-specific similarities between the image and descriptions. However, in our classname-free setup, it is unclear if plain averaging across the obtained classname-free descriptions and the single cls is appropriate. This is because the classname is probably the most important text representation of the class, whereas the classname-free descriptions rather have a supporting, distinguishing character. To address this challenge in our evaluation, a weighting factor $w_{cls}\in\mathbb{R}^{+}$ gets introduced to the vision-language ensemble:
@@
Weights of the classname-free descriptions are normalized to one in order to have the same relative weightings between classes with different amounts of assigned descriptions.
Nonetheless, the challenge remains how to find class-specific, classname-free descriptions that actually boost the classification accuracy, which we will discuss next.
@@
i.e., those $m$ descriptions whose averaged image-description similarity differences to $c$ are, on average, maximally large. In other words, these are descriptions that activate high for $a$ but not high for any $a^{\prime}\in A\setminus a$ on average. Because these $m$ descriptions are selected without a prepended class name cls, they can serve as classname-free language representations of class $c$.
@@
Better interpretability
Our proposed method achieves better interpretability by offering these four characteristics:
@@
DCLIP LLM assignments, global eval 13 61.99 55.09 51.79 43.31 39.91 43.09 62.97
DCLIP LLM assignments, local top-k eval 13 61.99 55.06 51.83 43.29 39.87 43.09 62.86
@@
Random Characters Ours 5 62.18 55.22 52.31 40.82 40.44 44.73 66.12
Contrastive LLM assignments 40 62.03 54.88 52.24 46.97 40.37 44.41 62.90
@@
Table 1: Image classification in classname-free setup. Our method consistently produces the highest accuracies in this setting. To ensure a fair comparison, we use the best performing $w_{cls}$ of the respective assignment.
@@ 4.1 Implementation Details and Datasets
Implementation details. We use CLIP (Radford et al. 2021) as the base Vision-Language Model (VLM) for our approach. Unless stated otherwise, the backbone for CLIP is the ViT-B/32 (Vaswani et al. 2017; Dosovitskiy et al. 2020). We randomly sample a subset from each dataset’s standard training split to obtain the lookup similarity table S (details see Section A.10). Our empirical tests confirmed that this sampling process does not significantly impact performance. The Large-Language Model (LLM) generated descriptors are sourced directly from DCLIP (Menon and Vondrick 2022) or generated using the contrastive prompting method with gpt-3.5-turbo-1106 and Llama-3-70b-chat-hf via APIs.
@@
Table 2: Performance in classname-free setup with $w_{cls}=0$. Our descriptions are robust enough and perform decently, even if the classname text $D_{cls}$ is weighted by $w_{cls}=0$. LLM assignments give a considerably worse performance in this scenario. Randomly assigned descriptions fail to provide reasonable guidance. Llama3-70B with Contrastive prompt is used as source pool $\mathcal{P}$.
@@
Table 3: Image classification with classname included in the descriptions. An ablation of ambiguous context size $k$ can be found in Section A.6.
LLMs ImageNet ImageNetV2 CUB200 EuroSAT Places365 DTD Flowers102
Contrastive Llama3-70B 61.68 54.64 53.49 36.33 41.33 79.87 64.45
w/ Our Selection 63.30 55.24 56.27 58.57 43.65 48.09 68.61
Contrastive GPT3.5 61.65 54.69 53.83 35.38 40.79 44.10 66.29
w/ Our Selection 63.15 55.09 56.21 44.43 43.64 80.78 68.48
@@ 4.2 Experimental Results
We evaluate the quality of the classname-free description assignments selected by our method in the classname-free evaluation setup (cf.  Section 3.2). Examples of selected descriptions can be found in  Section A.11. Performance of our algorithm across 7 classification benchmarks are shown in Figure 3, highlighting how varying $w_{cls}$ impacts Top-$1$ accuracy. The non-ensembled CLIP baseline performance, independent of $w_{cls}$, is also included for reference. Our selected assignments consistently outperform the DCLIP LLM assignments. Notably, for the EuroSAT, Flowers102, CUB200, DTD, and Places datasets, optimal performance occurs when $w_{cls}$ is low ($[0,10]$), emphasizing the importance of classname-free descriptions while exceeding the baseline performance by up to $9\%$ and the LLM performance by up to $8\%$. However, the LLM-assigned descriptions cannot produce performance gains in the classname-free scenario that comes close to our selections.
Further increasing $w_{cls}$ and thereby weighting the single classname-included description higher, reduces accuracy, showing that overly prioritizing the classname diminishes the benefits of our classname-free descriptions. Interestingly, the smaller gain for ImageNet (0.50.5~{}0.50.5pp.) also corresponds to a lower bump for low $w_{cls}$ in the plot. This may be due to the noisier backgrounds of this dataset, which hinders selecting generally valid descriptions.
@@
To confirm that our gains are not driven solely by the image-wise top-$k$ neighborhood, we also evaluate DCLIP LLM assignments in the local top-$k$ context, which shows no significant improvement. This suggests that our approach succeeds by the selection procedure within the top-$k$ neighborhood rather than the search space restriction alone. Importantly, these gains are independent of a specific description pool $\mathcal{P}$ as they hold for a contrastive prompting pool, too.
To our best knowledge, no prior work has explored a comparable classname-free evaluation setup to determine the true distinctiveness of assigned descriptions $d^{c-}_{0},...,d^{c-}_{m}$ in combination with a classname prompt $d^{c+}$. However, some works use methods like trainable bottleneck classifiers  (Yang et al. 2023; Yan et al. 2023) or trainable embeddings (Zang et al. 2024), which can be considered ”classname-free.” Despite this similarity, they are too different to compare against (detailed discussion in Section A.9).
@@
We evaluate our chosen descriptions in a conventional setup where classnames are included in all descriptions, as shown in Table 3. Our method performs well on datasets where a higher bump in performance is observed with low $w_{cls}$, i.e. high relative description weights in Figure 3. This happens when the selected description pool provides generalizable, diverse, and discriminative descriptions for the datasets. Our method outperforms DCLIP assignments in the DCLIP pool and outdoes WaffleCLip and FuDD on datasets CUB200, EuroSAT, Places365, DTD, and Flowers102. This remains true for $4$ of these datasets, even if only $5$ descriptions per class are used. In contrast, WaffleClip (Roth et al. 2023) uses $30$ text prompts per class, and FuDD (Esfandiarpoor and Bach 2024) uses an astonishing number of 1,842 descriptions per class.
On the other hand, for ImageNet and ImageNetV2, we can see a connection between suboptimal conventional performance and a much lower peak relative to baseline CLIP in the classname-free setting - indicating less distinctive power of our assignments. Mixed results in the conventional setup for ImageNet and ImageNetV2 imply it is challenging to find distinctive descriptions - at least within the currently used description pools $\mathcal{P}$. This difficulty may arise because random image contents, e.g. background objects, distort the description assignments. Our algorithm experiences a performance boost when using a $\mathcal{P}$ obtained through contrastive prompting, offering a richer pool of descriptions.
Overall, the results from the class name-containing scenario suggest that the added semantics of the discovered descriptions enhance performance—in addition to the class name ensembling used by other methods like WaffleClip, FuDD, and DCLIP.
@@
We evaluate the performance under a classname-free scenario in Table 2 without using any guiding classname information. In this case, random assignments don’t achieve any reasonable classification accuracy; LLM-assigned descriptions provide minor guidance. With our selected descriptions, however, we have achieved decent performance across all datasets - significantly surpassing the LLM assignments. This further supports the idea that descriptions assigned by an LLM are not distinctive enough. Instead, feedback from the embedding space is needed for distinctive assignments.
Combined with different LLMs.

### v2

@@ Abstract
Accurately describing images with text is a foundation of explainable AI. Vision-Language Models (VLMs) like CLIP have recently addressed this by aligning images and texts in a shared embedding space, expressing semantic similarities between vision and language embeddings. VLM classification can be improved with descriptions generated by Large Language Models (LLMs). However, it is difficult to determine the contribution of actual description semantics, as the performance gain may also stem from a semantic-agnostic ensembling effect, where multiple modified text prompts act as a noisy test-time augmentation for the original one. We propose an alternative evaluation scenario to decide if a performance boost of LLM-generated descriptions is caused by such a noise augmentation effect or rather by genuine description semantics. The proposed scenario avoids noisy test-time augmentation and ensures that genuine, distinctive descriptions cause the performance boost. Furthermore, we propose a training-free method for selecting discriminative descriptions that work independently of classname-ensembling effects. Our approach identifies descriptions that effectively differentiate classes within a local CLIP label neighborhood, improving classification accuracy across seven datasets. Additionally, we provide insights into the explainability of description-based image classification with VLMs.
@@
Extended Versionhttps://arxiv.org/abs/2412.11917
Figure 1: Are the extra semantics provided by LLM truly useful? Our method first identifies candidate labels using only the class name. We then filter out descriptions that may seem logical but do not differentiate the group, e.g. ambiguous, overly generic, or noisy descriptions. This refinement ensures that the remaining descriptions provide distinctive vision-language cues within the local candidate neighborhood, offering more specificity than the class name alone can capture.
@@ 1 Introduction
Human visual recognition is closely related to verbal reasoning, as it often relies on the ability to express visual information in words (Zhao et al. 2022; Shtedritski, Rupprecht, and Vedaldi 2023). However, a neural network usually does not exhibit this property, making its explainability a significant concern for the machine learning community. Some studies (Zhang et al. 2024; Hakimov and Schlangen 2023) aim to connect visual cues and textual descriptions, but these usually require extensive human subject analysis and highly specific datasets with annotations  (Young et al. 2014), which are expansive to obtain (Lin et al. 2014).
@@
Vision-Language Models (VLM) (Radford et al. 2021; Jia et al. 2021) tackle this issue by training neural networks to link images and their textual descriptions within a shared embedding space. This enhances the correlation between visual and textual details. VLMs can be applied to zero-shot image classification by passing an image through the VLM’s image encoder and prompting the text encoder with hand-crafted inputs like “a photo of a [classname]” (Radford et al. 2021). Recent work(Menon and Vondrick 2022) extends this approach by incorporating additional descriptions generated by Large Language Models (LLMs) for each class name. LLMs like GPT-3 (Brown et al. 2020; Ouyang et al. 2022) or Llama (Touvron et al. 2023), trained on extensive text corpora, are intended to provide richer semantics, enhancing VLM classification.
@@
However, Does VLM classification truly benefit from LLM-generated description semantics? This work explores this core question, as LLM-generated descriptions present several challenges. For example, descriptions can overlap for similar classes — such as parrots and sparrows — both described as having feathers, which is not a distinguishing feature. Moreover, while supplying the model with as many LLM-generated descriptions as possible may seem advantageous, it results in excessively lengthy collections. This complicates understanding the contribution of each description to the final decision.
@@
Another problem is the structured noise ensembling phenomenon (Roth et al. 2023): LLM-generated descriptions can be replaced with high-level concepts and random characters (such as “Baklava”, “a food that is 34mfqr5”) while still improving the model performance. These slightly modified duplications act as test-time augmentation for the original prompt, resulting in an averaged robust output. This raises the question whether the improvement is due to additional semantics of the LLM-generated descriptions or to the ensemble effect of the noise augmentation.
@@
Given these challenges, a proposed model should meet three criteria: 1) As humans who describe with a limited set of descriptions, the model should also operate with a manageable number of text descriptions. 2) These descriptions should be semantically meaningful. 3) The model should be resilient to noise ensembling. To address the issue of noise ensembling, we constrain the model to use only textual descriptions that do not contain the classname, i.e. classname-free descriptions.
@@ 3 Method
First, we introduce the conventional task formulation for image classification using Vision-Language Models. We then present our unique approach to this task to enable explainability. Finally, we propose a specific solution to enhance the results further.
@@ 3.1 Background
where $e(\cdot)$ is the image or language embedding. Another way to obtain vision-language scores is via ensembling. The motivation for ensembling can be derived from how a human describes an object. For example, when describing an apple, we can describe it as a “green stuff”, “a round object”, or “fruit of the same size as an orange”.
In this case, there is no single text representation $t_{c}$ for a class $c$ but a set of language representations $\mathcal{D}(c)$ where the ensembling happens over the elements of $\mathcal{D}(c)$:
@@
Figure 2: In the conventional setup (left), using CLIP with LLM-assigned class descriptions or even random strings can sometimes result in performance gains due to the added semantics or the smoothing ensemble effect. However, when the classname is removed, i.e. under the proposed classname-free setup (right), these descriptions will fail to perform well, as only meaningful descriptions w.r.t. the class are useful. In contrast, random strings or non-informative descriptions bring no gain.
@@ 3.2 Our Approach
In the conventional setup (Menon and Vondrick 2022; Pratt et al. 2023; Esfandiarpoor and Bach 2024), performance gains may result from the noise augmentation of the class name text [cls] embedding through its various combinations with [LLM-generated category], [LLM-generated descriptions], and even random strings. While random strings should not contribute extra semantics and are likely embedded far away from [cls], this can sometimes apply to LLM-assigned text due to the vocabulary discrepancy between VLMs and LLMs. Another cause may also be that the images do not exhibit the described property. Despite this, such combinations can still perform well, although the assigned descriptions are not semantically correct.
To better investigate whether the improved performance stems from semantic enrichment or the ensemble effect, we propose an approach where, out of all elements in $\mathcal{D}(c)$, exactly one element should contain the class name $c$. The remaining elements must contain textual descriptions without the class name. This set of descriptions then becomes:
@@
As discussed previously, the language-ensembling VLM method is evaluated under the conventional scenario by averaging the aggregated class-specific similarities between the images and class-specific descriptions.
However, in our classname-free setup, it is unclear if plain averaging across the obtained classname-free descriptions and the single cls is appropriate. This is because the classname is probably the most important text representation of the class, whereas the classname-free descriptions rather have a supporting, distinguishing character. To address this challenge in our evaluation, a weighting factor $w_{cls}\in\mathbb{R}^{+}$ gets introduced to the vision-language ensemble:
@@
Weights of the classname-free descriptions are normalized to one to have the same relative weightings between classes with different amounts of assigned descriptions. Nonetheless, the challenge remains how to find class-specific, classname-free descriptions that actually improve the classification accuracy. This we shall discuss next.
@@
i.e., those $m$ descriptions whose averaged image-description similarity differences to $c$ are, on average, maximally large. In other words, these descriptions activate highly for $a$ but not highly for any $a^{\prime}\in A\setminus a$ on average. Because these $m$ descriptions are selected without a prepended class name cls, they can serve as classname-free language representations of class $c$.
@@
Better explainability
Our proposed method achieves better explainability by offering these four characteristics:
@@
DCLIP LLM (global eval) 13 61.99 55.09 51.79 43.31 39.91 43.09 62.97
DCLIP LLM (local-k eval) 13 61.99 55.06 51.83 43.29 39.87 43.09 62.86
@@
Random Ours 5 62.18 55.22 52.31 40.82 40.44 44.73 66.12
Contrastive LLM 40 62.03 54.88 52.24 46.97 40.37 44.41 62.90
@@
Table 1: Image classification in classname-free setup with different assignments and pools. Our method consistently produces the highest accuracies in this setting. We use the best-performing $w_{cls}$ of the respective assignment to ensure a fair comparison.
@@ 4.1 Implementation Details and Datasets
Implementation details. We use CLIP (Radford et al. 2021) as the base Vision-Language Model (VLM) for our approach. Unless stated otherwise, the backbone for CLIP is the ViT-B/32 (Vaswani et al. 2017; Dosovitskiy et al. 2020). We randomly sample a subset from each dataset’s standard training split to obtain the lookup similarity table S (details see Section A.9). Our empirical tests confirmed that this sampling process does not significantly impact performance. The Large-Language Model (LLM) generated descriptions are sourced directly from DCLIP (Menon and Vondrick 2022) or generated using the contrastive prompting method with gpt-3.5-turbo-1106 and Llama-3-70b-chat-hf via APIs.
@@
Table 2: Performance in classname-free setup with $w_{cls}=0$. Our descriptions are robust and perform well, even if the classname text $D_{cls}$ is weighted by $w_{cls}=0$. LLM assignments give a considerably worse performance in this scenario. Randomly assigned descriptions fail to provide reasonable guidance. Llama3-70B with Contrastive Prompting is used as source pool $\mathcal{P}$. Ambiguous context size $k=3$. For sample sizes $n$ see Section A.9.
@@
Table 3: Image classification with classname included in the descriptions. Ambiguous context size $k=3$. For sample sizes $n$ see Section A.9. An ablation of ambiguous context size $k$ can be found in Section A.5.
@@ 4.2 Experimental Results
We evaluate the quality of the classname-free description assignments selected by our method in the classname-free evaluation setup (cf.  Section 3.2). Examples of selected descriptions can be found in  Section A.12. Performance of our algorithm across $7$ classification benchmarks is shown in Figure 3, highlighting how varying $w_{cls}$ impacts top-$1$ accuracy. The non-ensembled CLIP baseline performance, independent of $w_{cls}$, is also included for reference. Our selected assignments consistently outperform the DCLIP LLM assignments. Notably, for the EuroSAT, Flowers102, CUB200, DTD, and Places datasets, optimal performance occurs when $w_{cls}$ is low ($[0,10]$), emphasizing the importance of classname-free descriptions while exceeding the baseline performance by up to $9\%$ and the LLM performance by up to $8\%$. However, the LLM-assigned descriptions cannot produce performance gains in the classname-free scenario that comes close to our selections.
Further increasing $w_{cls}$ and thereby weighting the single classname-included description higher reduces accuracy, showing that overly prioritizing the classname diminishes the benefits of our classname-free descriptions.
Interestingly, the smaller gain for ImageNet ($\approx 0.5$pp.) also corresponds to a lower bump for low $w_{cls}$ in the plot. This may be due to the noisier backgrounds of this dataset, which hinders the selection of generally valid descriptions.
@@
To confirm that our gains are not driven solely by the image-wise top-$k$ neighborhood, we also evaluate DCLIP LLM assignments in the local top-$k$ context, which shows no significant improvement. This suggests that our approach succeeds by the selection procedure within the top-$k$ neighborhood rather than the search space restriction alone. Importantly, these gains are independent of a specific description pool $\mathcal{P}$ as they also hold for a contrastive prompting pool.
To our knowledge, no prior work has explored a comparable classname-free evaluation setup to determine the true distinctiveness of assigned descriptions $d^{c-}_{0},...,d^{c-}_{m}$ in combination with a classname prompt $d^{c+}$. However, some works use methods like trainable bottleneck classifiers  (Yang et al. 2023; Yan et al. 2023) or trainable embeddings (Zang et al. 2024), which can be considered ”classname-free.” Despite this similarity, they are too different to compare against (detailed discussion in Section A.8).
@@
We evaluate our chosen descriptions in a conventional setup where classnames are included in all descriptions, as shown in Table 3. Our method performs well on datasets where a higher performance bump is observed with low $w_{cls}$, i.e. high relative description weights in Figure 3. This happens when the selected description pool provides generalizable, diverse, and discriminative descriptions for the datasets. Our method outperforms DCLIP assignments in the DCLIP pool and outdoes WaffleCLip and FuDD on datasets CUB200, EuroSAT, Places365, DTD, and Flowers102. This remains true for $4$ of these datasets, even if only $5$ descriptions per class are used. In contrast, WaffleClip (Roth et al. 2023) uses $30$ text prompts per class, and FuDD (Esfandiarpoor and Bach 2024) uses an astonishing number of 1,842 descriptions per class.
On the other hand, for ImageNet and ImageNetV2, we can see a connection between suboptimal conventional performance and a much lower peak relative to baseline CLIP in the classname-free setting - indicating less distinctive power of our assignments. Mixed results in the conventional setup for ImageNet and ImageNetV2 imply it is challenging to find distinctive descriptions - at least within the currently used description pools $\mathcal{P}$. This difficulty may arise because random image contents, e.g., background objects, distort the description assignments. Our algorithm experiences a performance boost when using a $\mathcal{P}$ obtained through contrastive prompting, offering a richer pool of descriptions.
Overall, the results from the class name-containing scenario suggest that the added semantics of the discovered descriptions enhance the performance—in addition to the class name ensembling used by other methods like WaffleClip, FuDD, and DCLIP.
@@
We evaluate the performance under a classname-free scenario in Table 2 without any guiding classname information. In this case, random assignments don’t achieve any reasonable classification accuracy; LLM-assigned descriptions provide minor guidance. With our selected descriptions, however, we have achieved decent performance across all datasets - significantly surpassing the LLM assignments. This further supports the idea that descriptions assigned by an LLM are not distinctive enough. Instead, feedback from the embedding space is needed for distinctive assignments. Higher distinctiveness also shows in Section A.6 where classname ensembling is prohibited via a maxing-aggregation.